Question 2 View Policies Current Attempt in Progress A portion of the combined statement of income and retained earnings of WWaterway Inc. Income Statement TA
Answer #1

Answer:

Seminole Inc
Income Statement
For the year ended
Income before extra ordinary item $   15,260,000
Less: 6%Dividend on preferred stock $         318,000
Common stock income before extra ordinary item $   14,942,000
Less: Extra ordinary loss, net of tax $     1,330,000
Net income available for common stock holders $   13,612,000
Weighted Average number of shares outstanding
Date Outstanding Shares Fraction of years Weighted Shares
January 1 - April 1    7,520,000 3/12                    1,880,000
April 1 - December 31    8,540,000 9/12                    6,405,000
Weighted Average number of shares outstanding                    8,285,000
Earning Per Share
Income before extra ordinary loss $14,942,000/8,285,000 = $1.80
Extra ordinery loss $1,330,000/8,285,000 = $0.16
Net income / loss $13,612,000/8,285,000 = $1.64